# Directional expansiveness for $\mathbb{R}^{d}$-actions and for Penrose tilings

### v1

@@ 6.1. Wang tiles
There is a 1:1 correspondence between $Z_{q}\times[0,1)^{2}$ and $W_{q}$, in which $(z,{\boldsymbol{u}}^{\prime})$ with $z=(a_{\boldsymbol{n}})_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in Z_{q}$, and ${\boldsymbol{u}}^{\prime}\in[0,1)^{2}$ corresponds to $\{Q^{a_{\boldsymbol{n}}}+{\boldsymbol{n}}+{\boldsymbol{u}}^{\prime}\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in W_{q}$,
Now suppose $(Z,\mathbb{Z}^{2},S)$ is a $1$-step $\mathbb{Z}^{2}$ SFT, $Z\subseteq{\mathcal{A}}^{\mathbb{Z}^{2}}$, defined by the horizontal and vertical permitted words ${\mathcal{P}}_{0}$ and ${\mathcal{P}}_{1}$ (see Section 2.3). We say $(Z,\mathbb{Z}^{2},S)$ is colorable if for some finite sets ${\mathcal{C}}_{{\rm b},{\rm t}}$ and ${\mathcal{C}}_{{\rm l},{\rm r}}$, there are functions (i) $\xi_{\rm l},\xi_{\rm r}:{\mathcal{A}}\to{\mathcal{C}}_{{\rm l},{\rm r}}$ so that $aa^{\prime}\in{\mathcal{P}}_{0}$ if and only if $\xi_{\rm r}(a)=\xi_{\rm l}(a^{\prime})$, and (ii) $\xi_{\rm b},\xi_{\rm t}:{\mathcal{A}}\to{\mathcal{C}}_{{\rm b},{\rm t}}$ so that $(aa^{\prime})^{t}\in{\mathcal{P}}_{1}$ if and only if $\xi_{\rm b}(a)=\xi_{\rm t}(a^{\prime})$. Clearly, if $(Z_{q},\mathbb{Z}^{2},S)$ is the $1$-step $\mathbb{Z}^{2}$ SFT that comes from a Wang tile set $q$, then it is colorable. Conversely, suppose $(Z,\mathbb{Z}^{2},S)$ is a colorable $1$-step SPT. For each $a\in{\mathcal{A}}$ define the Wang tile $Q^{a}=(a,\xi_{\rm b}(a),\xi_{\rm t}(a),\xi_{\rm l}(a),\xi_{\rm r}(a))\in q$. Then $(Z_{q},\mathbb{Z}^{2},S)$ is identified with $(Z,\mathbb{Z}^{2},S)$ by the map that reads symbols. Note that our formulation of Wang tiles here allows for $Z_{q}={\mathcal{A}}^{\mathbb{Z}^{2}}$ to be a full shift.
@@ 6.2. Brief history
In terms of dynamics, aperiodicity for the Wang tile set $q$ means that $(W_{q},\mathbb{R}^{2},T)$ is free. Berger’s [2] apreiodic set $q$ of Wang tiles has #$\#(q)>20,\!000$, which led to a race to reduce this number (see [11] for a nice account). Recently, Jeandel and Rao [18] ended this race by constructing an aperiodic Wang tile set $q_{\rm jr}$ with #$\#(q_{\rm jr})=11$ and also showing #$\#(q_{\rm jr})=11$ is the smallest possible size.
@@ 6.3. Suspensions, compositions, deformations, and continuous orbit equivalence
A parallelogram Wang tiling dynamical system $(W_{Aq},\mathbb{R}^{2},T)$ is no longer the unit suspension of $(Z_{q},\mathbb{Z}^{2},S)$. However, we can recover it from $({\widetilde{Z_{q}}},\mathbb{R}^{2},{\widetilde{S}})$ as follows. For an $\mathbb{R}^{2}$-action $(X,\mathbb{R}^{2},T)$, and a matrix $A\in Gl(2,\mathbb{R})$, we define the composition action $(T\circ A)^{\boldsymbol{t}}x=T^{A{\boldsymbol{t}}}x$.
@@
For (1), we construct a model for ${\widetilde{S}}\circ A$ on $Z_{q}\times({\boldsymbol{f}}_{0}\wedge{\boldsymbol{f}}_{1})$. The conjugacy is the homeomorphism ${\rm id}\times A:Z_{q}\times[0,1)^{2}\to Z_{q}\times({\boldsymbol{f}}_{0}\wedge{\boldsymbol{f}}_{1})$. Part (2) follows from part (1), and the fact $(W_{q},\mathbb{R}^{2},T)$ and $(W_{q},\mathbb{R}^{2},T\circ A)$, (which is topologically conjugate to $(W_{Aq},\mathbb{R}^{2},T)$), both act on $W_{q}$, and have the same orbits. Part (3) follows from Propositions 3.12 and 3.9. ∎
@@
    Suppose $\widetilde{T}$ is the unit suspension of $T$. Then $\Sigma_{\widetilde{T}}=\pi^{-1}\Sigma_{T}$ where $\pi({\boldsymbol{s}}):=\{{\boldsymbol{s}}\}:\mathbb{R}^{2}\to\mathbb{T}^{2}$.
@@ 23 $\overline{2}\overline{3}4,4\overline{2},\ \overline{3}\overline{4},\overline{4}$ 3,1, 5,4 0,1, 3,2 9
Column 2 of Table 1 shows that there are nine distinct top and bottom code words, and for each of them we assign a color $c_{\rm b},c_{\rm t}\in{\mathcal{C}}_{{\rm b},{\rm t}}=\{0,1,\dots,8\}$. Similarly, there are nine distinct left and right code words assigned colors $c_{\rm l},c_{\rm r}\in{\mathcal{C}}_{{\rm l},{\rm r}}=\{0,1,\dots,8\}$. The colors are shown in Table 1, Column 3.
@@
Two tiles $Q^{a}$ and $Q^{a^{\prime}}$ can be vertically (or horizontally) adjacent in some Wang tiling in $W_{q_{\rm p}}$ if and only if the same is true for the corresponding Wang patches $w^{a}$ and $w^{a^{\prime}}$ for some $x\in\mathsf{X}$. The Penrose Wang tile dynamical system is denoted $(W_{q_{\rm p}},\mathbb{R}^{2},T)$, and the corresponding $\mathbb{Z}^{2}$ SFT is denoted $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$, where we write $(a_{\boldsymbol{n}})_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in Z_{q_{\rm p}}$.
@@
Let $\{Q^{a_{\boldsymbol{n}}}+{\boldsymbol{n}}+{\boldsymbol{u}}^{\prime}\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in W_{q_{\rm p}}$, ${\boldsymbol{u}}^{\prime}\in[0,1)^{2}$, $(a_{\boldsymbol{n}})_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in Z_{q_{\rm p}}$, be an arbitrary Penrose Wang tiling, ((((see Remark 6.1)))). Then there is a Penrose tiling $x\in\mathsf{X}$, unique up to translation, so that $w_{\boldsymbol{n}}(x)$ is type $w^{a_{\boldsymbol{n}}}$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. Conversely, for each Penrose tiling $x\in\mathsf{X}$ there is a Penrose Wang tiling $\{Q^{a_{\boldsymbol{n}}}+{\boldsymbol{n}}\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$, unique up to translation, so that $w_{\boldsymbol{n}}(x)$ is type $w^{a_{\boldsymbol{n}}}$ for each ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. In particular, $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ and $(\mathsf{X},\mathbb{R}^{2},T))$ and are continuous orbit equivalent (via a deformation), and $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ and $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$ are nonempty and free.
@@
Like $(W_{q_{\rm p}},\mathbb{R}^{2},T)$, the Penrose parallelogram Wang tile dynamical system $(W_{Aq_{\rm p}},\mathbb{R}^{2},T)$, nonempty, free and is Kakutanni equivalent to $(\mathsf{X},\mathbb{R}^{2},T)$. More of its properties are described in Theorem 6.11 below.
@@
These vectors can also be found by summing the edge vectors of the along the sides of the corresponding Wang patch $w_{\boldsymbol{n}}(x)$:
@@
Suppose $U_{\boldsymbol{n}}(x)$ is the unmarked tetragon Penrose tile corresponding to the Wang patch $w_{\boldsymbol{n}}(x)$ in some $x\in\mathsf{X}$. If the type of $w_{\boldsymbol{n}}(x)$ is $w^{a}$ (so that we think of $U_{\boldsymbol{n}}$ as type $U^{a}$). Then for each $s\in\{{\rm b},{\rm t}\}$ and $s^{\prime}\in\{{\rm l},{\rm r}\}$
@@
In particular ${\boldsymbol{b}}_{s}^{a}$ for $s\in\{{\rm b},{\rm t},{\rm l},{\rm r}\}$ depends only on the code word $\beta_{s}^{a}$, or equivalently, the color $c_{s}^{a}$ (Table1, Column 2).
@@
Figure 12. The 24 tetragon Penrose tiles $p_{4}:=\{M^{a}:a\in{\mathcal{A}}\}$ is shown in the rounded squares. Each edge is labeled with its color, $c_{\rm b},c_{\rm t}\in\{{\it 0},\dots\it 8\}$ and $c_{\rm l},c_{\rm r}\in\{{\it 0},\dots\it 8\}$, (Table 1, Column 3), with the subscript showing the side vector $v$ or $v^{\prime}$ (Table 2, Columns 2 and 4). The tile shape (the unmarked tetragon tile type $v$) is show in the circle below each tile. The 24 Penrose Wang tiles (or Penrose parallelogram Wang tiles) may be obtained by deforming each tetragon with the unit square or rhomb.
@@
We define marked tetragon Penrose tiles $p^{4}=\{M^{a},a\in{\mathcal{A}}\}$ as follows. For each $a\in{\mathcal{A}}$ (Table 1, Column 1) we define $M^{a}$ by decorating the corresponding unmarked tetragon Penrose tile $U^{v}$, where $v=v(a)$ is given in Table 1, Column 5, by coloring its edges using the Wang tile colors $c_{\rm b},c_{\rm t},c_{\rm l}$, and $c_{\rm r}$ (Table 1, Column 3). The matching rule $p_{4}^{(2)}$ is the requirement that the colors of the adjacent edges of two adjacent tiles must match.
Any two horizontal or vertical or edges that have the same color also have the same edge vector. Thus the adjacency condition is the same as for the Penrose Wang tiles. We call $(X_{p_{4}},\mathbb{R}^{2},T)$ the tetragon Penrose tiling dynamical system. We see that $(X_{p_{4}},\mathbb{R}^{2},T)$ is continuously orbit equivalent to $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ by a deformation, so in particular, $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ is nonempty and free.
The next two results summarize the dynamical properties of the four dynamical systems studied in this section. Theorem 6.11 discusses the tetragon Penrose tiling system $(X_{p_{4}},\mathbb{R}^{2},T)$ and Penrose parallelogram Wang tile system $(W_{Aq_{\rm p}},\mathbb{R}^{2},T)$. Corollary 6.12 discusses Raphael Robinson’s Penrose Wang tile system $(W_{q_{\rm p}},\mathbb{R}^{2},T)$, and the corresponding $\mathbb{Z}^{2}$ SFT $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$.
@@
The dynamical systems $(X_{p_{4}},\mathbb{R}^{2},T)$ and $(W_{Aq_{\rm p}},\mathbb{R}^{2},T)$ are both topologically conjugate ((((MLD in the first case)))) to the Penrose tiling dynamical system $(\mathsf{X},\mathbb{R}^{2},T)$. Each is strictly ergodic and almost automorphic, with $\Sigma_{T}=Z[e^{2\pi i/5}]\subseteq\mathbb{R}^{2}$, $\mathfrak{M}_{T}=\{1,2,10\}$, and $\mathbb{N}_{1}(T)=\{{\boldsymbol{v}}_{0}^{\perp},\dots,{\boldsymbol{v}}_{4}^{\perp}\}$.
@@
This follows from Theorem 6.11, Lemma 6.3, and Remark 6.4. However $\Sigma_{S}=(\mathbb{Z}[\alpha]/\mathbb{Z})\times(\mathbb{Z}[\alpha]/\mathbb{Z})$ is easier to see from the fact that $R_{\alpha}\otimes R_{\alpha}$ is the Kronecker factor of $S$. ∎
@@
In [16], the minimality of $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$ is stated as a conjecture, and an earlier preprint version of [16], which also also proves that $\mathbb{N}_{1}(T)$ is as stated above.
@@
It suffices to prove the topological conjugacies, and then everything else follows from Theorems 4.3 and 1.1. For $(X_{p_{4}},\mathbb{R}^{2},T)$ there is an obvious MLD topological conjugacy.

### v2

@@ 6.1. Wang tiles
There is a 1:1 correspondence between $Z_{q}\times[0,1)^{2}$ and $W_{q}$, in which $(z,{\boldsymbol{u}}^{\prime})$ with $z=(a_{\boldsymbol{n}})_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in Z_{q}$, and ${\boldsymbol{u}}^{\prime}\in[0,1)^{2}$ corresponds to $\{Q^{a_{\boldsymbol{n}}}+{\boldsymbol{n}}+{\boldsymbol{u}}^{\prime}\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in W_{q}$.
Now suppose that $(Z,\mathbb{Z}^{2},S)$ is a $1$-step $\mathbb{Z}^{2}$ SFT, $Z\subseteq{\mathcal{A}}^{\mathbb{Z}^{2}}$, defined by the horizontal and vertical permitted words ${\mathcal{P}}_{0}$ and ${\mathcal{P}}_{1}$ (see Section 2.3). We say $(Z,\mathbb{Z}^{2},S)$ is colorable if for some finite sets ${\mathcal{C}}_{{\rm b},{\rm t}}$ and ${\mathcal{C}}_{{\rm l},{\rm r}}$, there are functions (i) $\xi_{\rm l},\xi_{\rm r}:{\mathcal{A}}\to{\mathcal{C}}_{{\rm l},{\rm r}}$ so that $aa^{\prime}\in{\mathcal{P}}_{0}$ if and only if $\xi_{\rm r}(a)=\xi_{\rm l}(a^{\prime})$, and (ii) $\xi_{\rm b},\xi_{\rm t}:{\mathcal{A}}\to{\mathcal{C}}_{{\rm b},{\rm t}}$ so that $(aa^{\prime})^{t}\in{\mathcal{P}}_{1}$ if and only if $\xi_{\rm b}(a)=\xi_{\rm t}(a^{\prime})$. Clearly, if $(Z_{q},\mathbb{Z}^{2},S)$ is the $1$-step $\mathbb{Z}^{2}$ SFT that comes from a Wang tile set $q$, then it is colorable. Conversely, suppose that $(Z,\mathbb{Z}^{2},S)$ is a colorable $1$-step SFT. For each $a\in{\mathcal{A}}$ define the Wang tile $Q^{a}:=(a,\xi_{\rm b}(a),\xi_{\rm t}(a),\xi_{\rm l}(a),\xi_{\rm r}(a))\in q$. Then $(Z_{q},\mathbb{Z}^{2},S)$ is identified with $(Z,\mathbb{Z}^{2},S)$ by the map that reads symbols. Note that our formulation of Wang tiles here allows for $Z_{q}={\mathcal{A}}^{\mathbb{Z}^{2}}$ to be a full shift.
@@ 6.2. Brief history
In terms of dynamics, aperiodicity for the Wang tile set $q$ means that $(W_{q},\mathbb{R}^{2},T)$ is free. Berger’s [2] apreiodic set $q$ of Wang tiles has #$\#(q)>20,\!000$, which led to a race to reduce this number (see [11] for a nice account). Recently, Jeandel and Rao [18] ended this race by constructing an aperiodic Wang tile set $q_{\rm jr}$ with #$\#(q_{\rm jr})=11$ and also showing that #$\#(q_{\rm jr})=11$ is the smallest possible size.
@@ 6.3. Suspensions, compositions, deformations, and continuous orbit equivalence
A parallelogram Wang tiling dynamical system $(W_{Aq},\mathbb{R}^{2},T)$ is no longer the unit suspension of $(Z_{q},\mathbb{Z}^{2},S)$. However, we can recover it from $({\widetilde{Z_{q}}},\mathbb{R}^{2},{\widetilde{S}})$ as follows: For an $\mathbb{R}^{2}$-action $(X,\mathbb{R}^{2},T)$, and a matrix $A\in Gl(2,\mathbb{R})$, we define the composition action by $(T\circ A)^{\boldsymbol{t}}x:=T^{A{\boldsymbol{t}}}x$.
@@
For (1), we construct a model for ${\widetilde{S}}\circ A$ on $Z_{q}\times({\boldsymbol{f}}_{0}\wedge{\boldsymbol{f}}_{1})$. The conjugacy is the homeomorphism ${\rm id}\times A:Z_{q}\times[0,1)^{2}\to Z_{q}\times({\boldsymbol{f}}_{0}\wedge{\boldsymbol{f}}_{1})$. Part (2) follows from part (1), and the fact $(W_{q},\mathbb{R}^{2},T)$ and $(W_{q},\mathbb{R}^{2},T\circ A)$, (which is topologically conjugate to $(W_{Aq},\mathbb{R}^{2},T)$), both act on $W_{q}$, and have the same orbits. Part (3) follows from Propositions 3.13 and 3.10. ∎
@@
    Suppose that $\widetilde{T}$ is the unit suspension of $T$. Then $\Sigma_{\widetilde{T}}=\pi^{-1}\Sigma_{T}$ where $\pi({\boldsymbol{s}}):=\{{\boldsymbol{s}}\}:\mathbb{R}^{2}\to\mathbb{T}^{2}$.
@@ 23 $\overline{2}\overline{3}4,4\overline{2},\ \overline{3}\overline{4},\overline{4}$ 3,1, 5,4 0,1, 3,2 9
Column 2 of Table 1 shows that there are nine distinct top and bottom code words, and for each of them we assign a color $c_{\rm b},c_{\rm t}\in{\mathcal{C}}_{{\rm b},{\rm t}}=\{0,1,\dots,8\}$. Similarly, there are nine distinct left and right code words each assigned a color $c_{\rm l},c_{\rm r}\in{\mathcal{C}}_{{\rm l},{\rm r}}=\{0,1,\dots,8\}$. The colors are shown in Table 1, Column 3.
@@
Two tiles $Q^{a}$ and $Q^{a^{\prime}}$ can be vertically (or horizontally) adjacent in some Wang tiling in $W_{q_{\rm p}}$ if and only if the same is true for the corresponding Wang patches $w^{a}$ and $w^{a^{\prime}}$ for some $x\in\mathsf{X}$. The Penrose Wang tile dynamical system is denoted by $(W_{q_{\rm p}},\mathbb{R}^{2},T)$, and the corresponding $\mathbb{Z}^{2}$ SFT is denoted by $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$, where we write $(a_{\boldsymbol{n}})_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in Z_{q_{\rm p}}$.
@@
Let $\{Q^{a_{\boldsymbol{n}}}+{\boldsymbol{n}}+{\boldsymbol{u}}^{\prime}\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in W_{q_{\rm p}}$, ${\boldsymbol{u}}^{\prime}\in[0,1)^{2}$, $(a_{\boldsymbol{n}})_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}\in Z_{q_{\rm p}}$, be an arbitrary Penrose Wang tiling, ((((see Remark 6.1)))). Then there is a Penrose tiling $x\in\mathsf{X}$, unique up to translation, so that $w_{\boldsymbol{n}}(x)$ is type $w^{a_{\boldsymbol{n}}}$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. Conversely, for each Penrose tiling $x\in\mathsf{X}$ there is a Penrose Wang tiling $\{Q^{a_{\boldsymbol{n}}}+{\boldsymbol{n}}\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$, unique up to translation, so that $w_{\boldsymbol{n}}(x)$ is type $w^{a_{\boldsymbol{n}}}$ for each ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. In particular, $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ and $(\mathsf{X},\mathbb{R}^{2},T)$ are continuously orbit equivalent (via a deformation), and $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ and $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$ are nonempty and free.
@@
Like $(W_{q_{\rm p}},\mathbb{R}^{2},T)$, the Penrose parallelogram Wang tile dynamical system $(W_{Aq_{\rm p}},\mathbb{R}^{2},T)$ is nonempty, free, and is Kakutanni equivalent to $(\mathsf{X},\mathbb{R}^{2},T)$. More of its properties are described in Theorem 6.11 below.
@@
These vectors can also be found by summing the edge vectors along the sides of the corresponding Wang patch $w_{\boldsymbol{n}}(x)$:
@@
Suppose that $U_{\boldsymbol{n}}(x)$ is the unmarked tetragon Penrose tile corresponding to the Wang patch $w_{\boldsymbol{n}}(x)$ in some $x\in\mathsf{X}$ ((((we think of $U_{\boldsymbol{n}}$ as type $U^{a}$ if the type of $w_{\boldsymbol{n}}(x)$ is $w^{a}$)))). Then for each $s\in\{{\rm b},{\rm t}\}$ and $s^{\prime}\in\{{\rm l},{\rm r}\}$,
@@
In particular, ${\boldsymbol{b}}_{s}^{a}$ for $s\in\{{\rm b},{\rm t},{\rm l},{\rm r}\}$ depends only on the code word $\beta_{s}^{a}$, or equivalently, the color $c_{s}^{a}$ (Table 1, Column 2).
@@
Figure 13. The 24 tetragon Penrose tiles $p_{4}:=\{M^{a}:a\in{\mathcal{A}}\}$, where $a$ is shown in the rounded squares. Each edge is labeled with its color, $c_{\rm b},c_{\rm t}\in\{{\it 0},\dots\it 8\}$ and $c_{\rm l},c_{\rm r}\in\{{\it 0},\dots\it 8\}$, (Table 1, Column 3), with the subscript showing the side vector $v$ or $v^{\prime}$ (Table 2, Columns 2 and 4). The tile shape (the unmarked tetragon tile type $v$) is shown in the circle below each tile. The 24 Penrose Wang tiles (or Penrose parallelogram Wang tiles) may be obtained by deforming each tetragon with the unit square or rhomb.
@@
We define marked tetragon Penrose tiles $m_{4}=\{M^{a},a\in{\mathcal{A}}\}$ as follows. For each $a\in{\mathcal{A}}$ (Table 1, Column 1) we define $M^{a}$ by decorating the corresponding unmarked tetragon Penrose tile $U^{v}$, where $v=v(a)$ is given in Table 1, Column 5, by coloring its edges using the Wang tile colors $c_{\rm b},c_{\rm t},c_{\rm l}$, and $c_{\rm r}$ (Table 1, Column 3). The matching rule $m_{4}^{(2)}$ is the requirement that the colors of the adjacent edges of two adjacent tiles must match.
Any two horizontal or vertical edges that have the same color also have the same edge vector. Thus the adjacency condition is the same as for the Penrose Wang tiles. We call $(X_{m_{4}},\mathbb{R}^{2},T)$ the tetragon Penrose tiling dynamical system. We see that $(X_{m_{4}},\mathbb{R}^{2},T)$ is continuously orbit equivalent to $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ by a deformation, so in particular, $(W_{q_{\rm p}},\mathbb{R}^{2},T)$ is nonempty and free.
The next two results summarize the dynamical properties of the four dynamical systems studied in this section. Theorem 6.11 discusses the tetragon Penrose tiling system $(X_{m_{4}},\mathbb{R}^{2},T)$ and Penrose parallelogram Wang tile system $(W_{Aq_{\rm p}},\mathbb{R}^{2},T)$. Corollary 6.12 discusses Raphael Robinson’s Penrose Wang tile system $(W_{q_{\rm p}},\mathbb{R}^{2},T)$, and the corresponding $\mathbb{Z}^{2}$ SFT $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$.
@@
The dynamical systems $(X_{m_{4}},\mathbb{R}^{2},T)$ and $(W_{Aq_{\rm p}},\mathbb{R}^{2},T)$ are both topologically conjugate ((((MLD in the first case)))) to the Penrose tiling dynamical system $(\mathsf{X},\mathbb{R}^{2},T)$. Each is strictly ergodic and almost automorphic, with $\Sigma_{T}=Z[e^{2\pi i/5}]\subseteq\mathbb{R}^{2}$, $\mathfrak{M}_{T}=\{1,2,10\}$, and $\mathbb{N}_{1}(T)=\{{\boldsymbol{v}}_{0}^{\perp},\dots,{\boldsymbol{v}}_{4}^{\perp}\}$.
@@
This follows from Theorem 6.11, Lemma 6.3, and Remark 6.4. However, $\Sigma_{S}=(\mathbb{Z}[\alpha]/\mathbb{Z})\times(\mathbb{Z}[\alpha]/\mathbb{Z})$ is easier to see from the fact that $R_{\alpha}\otimes R_{\alpha}$ is the Kronecker factor of $S$. ∎
@@
In [16], the minimality of $(Z_{q_{\rm p}},\mathbb{Z}^{2},S)$ is stated as a conjecture, and an earlier preprint version of [16]. The preprint [16] also proves that $\mathbb{N}_{1}(T)$ is as stated above.
@@
It suffices to prove the topological conjugacies, and then everything else follows from Theorems 4.3 and 1.1. For $(X_{m_{4}},\mathbb{R}^{2},T)$ there is an obvious MLD topological conjugacy.